How many liters of a 60% acid solution must be mixed with a 75% acid solution to get 20 liters of 72% solution?
Answer by checkley71(8403)

.75x+.60(20-x)=.72*20
.75x+12-.60x=14.4
.15x=14.4-12
.15x=2.4
x=2.4/.15
x=16 liters of 75 % soluton is used.
20-16=4 liters of 60% solution is used.
proof
.75*16+.6*4=14.4
12+2.4=14.4
14.4=14.4